OpenStudy (kimberly_pr):

Solve for v 3v + 9 - 11v = -23

OpenStudy (kimberly_pr):

This is my last question, yayyyy

OpenStudy (kimberly_pr):

@Will.H

OpenStudy (will.h):

any thoughts?

OpenStudy (will.h):

HINT: you should simplify alike terms such that 3v and -11v 3v - 11v

OpenStudy (kimberly_pr):

I always get confused when there's 2 variables in an equation instead of one :/ so both v values have to be the same right?

OpenStudy (kimberly_pr):

Would it simplify to 9- 8v = -23?

OpenStudy (will.h):

no no it is 3v - 11v = ?v HINT: you can draw a numbers line like start counting

OpenStudy (will.h):

or you can set it like this -11v + 3v = ?v -11 add to it 3 positive numbers it would equal -8v

OpenStudy (will.h):

make any sense?

OpenStudy (kimberly_pr):

Sort of... But wouldn't I still end up with -8v?

OpenStudy (will.h):

TERRIFIC that's right. so we simplified the terms with variables and now we have the expression as -8v + 9 = -23

OpenStudy (will.h):

can you solve it from here?

OpenStudy (kimberly_pr):

The answer is 4!! :)

OpenStudy (will.h):

Great work ;)

OpenStudy (kimberly_pr):

Thabk you for all your help today :) I fanned you. And I'll make sure to tag you in the furutre when I need more help @Will.H

OpenStudy (kimberly_pr):

*Thank

OpenStudy (will.h):

you're welcome happy to help :) have a good day

OpenStudy (kimberly_pr):

*future